# A Survey of Artificial Intelligence in Gait-Based Neurodegenerative Disease Diagnosis

### v2

@@ 4.7 Amyotrophic Lateral Sclerosis (ALS)
In [104] (Table 7, ID = S-1), the study focuses on analyzing gait variability in patients with amyotrophic lateral sclerosis (ALS) and extracting two key features, standard deviation statistics and permutation entropy. These features were inputted into a support vector machine classifier to classify ALS patients and healthy controls. The results demonstrated high accuracy (92.86%) in distinguishing ALS patients from controls, indicating the potential of gait variability analysis as a diagnostic or monitoring tool for ALS. [224] (ID = S-2) leverages one-minute gait series and parameters to classify between ALS patients and healthy subjects using machine learning models (SVM, KNN, DT) and achieves up to 96.6% accuracy with DT. [225] (ID = S-3) extracts metrics of fluctuation magnitude and fluctuation dynamics from gait time series, and leverages five classic machine learning models (SVM, KNN, NB, LDA, DT) to classify ALS and healthy subjects. Both SVM and KNN achieves the best performance with 89.7% accuracy on the GaitNDD database.
@@ 5.2 Scarcity and Imbalance of Data
The medical data including gait data collected for NDs diagnosis are highly limited, and in the selected 164 articles only several datasets are publicly available: The Gait in Neurodegenerative Disease Database and Gait in Parkinson’s Disease from Physionet database [234] are two most frequently-used used datasets; Other datasets include SDUGait [235], Floodlight PoC dataset [236], UK Biobank Dataset [237], Ga, Ju, Si Groups [238, 239, 240]. However, these datasets are limited in both sample sizes and NDs types. Most of them contain less than 100 subjects (see Sec. 3.2.6), and the number of NDs types is typically less than 3, while PD gait data are much more than other NDs in terms of quantity and quality ($i.e.,$ unbalanced data size of different diseases). This could limit the learning performance of AI models especially deep learning models, which inherently require large-scale and class-balanced data to achieve more reliable prediction.
@@ 5.3 Lack of Multi-Source/Modal Gait Data
Therefore, an important future direction is to collect gait data from multiple sources and modalities. First, it is feasible to combine multiple devices as different sources of same-type data ($e.g.,$ multi-view data with differently-positioned cameras), where all devices/sources can compensate each other to better reconstruct or depict the whole pattern with less information loss [242]. Second, combining different modalities such as RGB images, depth images, 3D skeletons, and force data (vGRF) is valuable as they can provide gait information from different dimensions ($e.g.,$ appearances, poses) to better predict NDs [77].
@@ 5.4 Challenges in AI Model Design
To address the aforementioned challenges, it is essential to improve not only the accuracy of an AI model but also its generalizability, efficiency, and interpretability. A potential future direction is to exploit more advanced deep learning models and larger-scale gait datasets with different scenario settings to learn both domain-specific ($e.g.,$ AD-specific) and domain-general ($i.e.,$ NDs-shared) gait features, so as to improve the model accuracy and enable it to be transferred/generalized to other scenarios. For efficiency, it is feasible to employ lighter network architecture ($e.g.,$ MobileNets [250]) or model compression techniques ($e.g.,$ knowledge distillation [251], network pruning [252]) to reduce the model size and computational complexity but retaining similar model accuracy. Another promising direction is to adopt smaller gait data such as the emerging 3D skeleton data, which can concisely represent human poses with key body joints and typically require lightweight AI models to learn discriminative gait patterns [253]. To improve model interpretability, different types of human-friendly explanation including text description and feature visualization can be considered. We can integrate model-specific explanation mechanisms into different AI architectures, $e.g.,$ visual class activation maps (CAM) [254] for CNN, knowledge graphs [255] for graph neural networks (GNN). Inspired by the recent success of ChatGPT, it can be used to provide insightful description for the importance of features if we transform gait features into text format. Moreover, ChatGPT can also act as an agent to lead and improve the AI model learning [256], in which we can add prompts to generate interpretable results. As there exist key body parts and poses for NDs diagnosis, we can try to disentangle the gait abnormality related semantics from normal motion semantics ($e.g.,$ daily actions) to identify the most important or unique component of NDs patients’ features. Simultaneously improving the accuracy, generalizability, efficiency, and interpretability, it is hopefully to obtain a more reliable and practical AI model for gait-based NDs diagnosis. Meanwhile, a unified evaluation protocol, which synergizes different benchmark datasets, validation settings, and evaluation metrics, should be further devised to fairly compare different AI models to facilitate the development of this area.
@@ 6 Conclusion
In this paper, we provide an overview of the recent advancements of Artificial Intelligence (AI) models for diagnosing neurodegenerative diseases (NDs) through human gait. Our survey showcases a general process of gait-based NDs diagnosis using AI with focus on five representative NDs: PD, AD, HD, ALS, MS, and systematically reviews all existing studies with an elaboration on their used gait data, AI models, and overall performance. A novel quality evaluation criterion is also devised to quantitatively assess the quality of surveyed studies. We outline and discuss the current challenges, potential solutions, and promising future directions in this field. Furthermore, we propose a novel research vision on the utilization of emerging 3D skeleton data for human gait representation and AI model development for more efficient NDs diagnosis.

### v3

@@ 4.7 Amyotrophic Lateral Sclerosis (ALS)
In the work by Xia et al. [115] (Table 7, ID = S-1), they focused on analyzing gait variability in patients with ALS and extracted two key features, standard deviation statistics and permutation entropy. These features were inputted into a support vector machine classifier to classify ALS patients and healthy controls. The results demonstrated high accuracy (92.86%) in distinguishing ALS patients from controls, indicating the potential of gait variability analysis as a diagnostic or monitoring tool for ALS. Felix et al. [236] (ID = S-2) leveraged one-minute gait series and parameters to classify between ALS patients and healthy subjects using machine learning models (SVM, KNN, DT) and achieved up to 96.6% accuracy with DT. Felix et al. [237] (ID = S-3), in their another work, extracted metrics of fluctuation magnitude and fluctuation dynamics from gait time series, and leveraged five classic ML models (SVM, KNN, NB, LDA, DT) to classify ALS and healthy subjects. Both SVM and KNN achieved the best performance with a 89.7% accuracy on the GaitNDD database.
5 Comparison with Existing Surveys
To highlight the distinct contributions of our survey, we provide a brief comparison with the latest representative surveys or reviews on similar topics [39, 246, 247] in this section.
  1. 1.
    Comparison with the review by Keserwani et al. [246]: First, it only surveyed the PD and its related AI algorithms, while our survey simultaneously focused on AI techniques for the diagnosis of five typical NDs, including PD, AD, ALS, HD, and MS. Our survey is more comprehensive than this review by including a larger number of representative studies (32 more studies) and cited papers (85 more papers). Second, the focused main modality and its taxonomy are significantly different. The authors concentrated on the medical data such as neuroimaging, and regarded gait as a single modality without further categorizing its data types. In practice, gait can also be captured from different modalities such as RGB videos and force signals ($e.g.,$ vGRF), which requires a finer data classification. In our survey, with a focus on gait-based diagnosis, we present a more systematic taxonomy of gait data in terms of different sensor modalities, vision modalities, their combined modalities, and others (see Sec. 2.4). Moreover, this review only briefly mentioned several challenges without detailing their potential solutions and future directions, while our survey provided a more comprehensive discussion on different challenges with an elaboration on the potential future solution for each challenge. We further offer a novel research vision on the application of 3D skeleton data (see Sec. 7).
  2. 2.
    Comparison with the review by Gupta et al. [39]: (1) Firstly, this review only concerned with Parkinson’s Disease (PD) related research, while our review covered the 5 most prevalent neurodegenerative diseases (NDs), including PD, Alzheimer’s disease (AD), Amyotrophic Lateral Sclerosis (ALS), Huntington’s Disease (HD), and Multiple Sclerosis (MS). (2) Secondly, the focus of the data in this review and ours are distinctly different as ours focused entirely on works that harness different types of gait data. For instance, sensor-based (e.g., Inertial Measurement Unit (IMU), force-sensitive resistors), vision-based (e.g., 3D motion capture system), and multi-modal (e.g., Microsoft Kinect) gait data. However, the compared review [39] focused on works that explored various PD data modalities, such as biomarkers, speech recording, handwriting data. (3) In addition, the authors reviewed those works that target at different purposes such as PD diagnosis, treatment, management, identification of novel biomarkers in the progression of disease, and using AI to advance neurosurgical process and drug discovery, while our focus was solely on NDs’ diagnosis. (4) The review also discussed the potential role of the metaverse, the Internet of Things, and electronic health records in the effective management of PD to improve the quality of life. However, it lacks a discussion regarding current biggest challenges in current field of interest such as data scarcity, multi-modal integration, and the design of AI models which are detailed out in our review. What’s more, we also proposed harnessing novel 3D skeleton-based gait data to further advance the development of the field by addressing the state-of-the-art approaches related.
  3. 3.
    Comparison with the review [247]: (1) Compared to review [247] that focuses solely on Parkinson’s Disease (PD) and deep learning (DL) techniques through bibliometric analysis and method review, our study is broader and deeper in scope, methodology, and analysis. First, while the survey is limited to PD, we cover five major neurodegenerative diseases (NDs), including Alzheimer’s Disease (AD), amyotrophic lateral sclerosis (ALS), Huntington’s Disease (HD), and Multiple Sclerosis (MS), providing a more comprehensive perspective. (2) In terms of diagnostic methods, the review [247] emphasizes various PD-related indicators (e.g., biomarkers, handwriting, and speech analysis), whereas we focus on gait data as a key motor symptom, introducing a systematic taxonomy of gait data types (e.g., sensor modalities, vision modalities, and their combinations). (3) While the review [247] utilizes VOSviewer for bibliometric analysis, we not only review more studies (164) but also propose an innovative quality evaluation criterion to quantify the quality of existing research, adding depth to our analysis. (4) The review [247] highlights research gaps in incremental learning and big data analysis without offering detailed solutions, whereas we comprehensively discuss challenges such as data scarcity, multi-modal integration, and AI model design, proposing potential solutions and envisioning the application of 3D skeleton data and the development of efficient AI models. Through broader disease coverage, a stronger focus on gait analysis, and a detailed discussion of challenges and solutions, our survey provides systematic and forward-looking guidance for AI research in ND diagnosis.
6 Representative Datasets and Methods
6.1 Public Datasets
According to our survey, many studies manually collect their own gait data without publishing them as a public benchmark. Here we introduce several most commonly used gait datasets.
  • Gait in Neurodegenerative Disease Database [248]. This database includes force-based records from patients with Parkinson’s disease (15 persons), Huntington’s disease (20 persons), and amyotrophic lateral sclerosis (13 persons). Additionally, records from 16 healthy control subjects are provided. The raw data were collected using force-sensitive resistors, which produce an output roughly proportional to the force exerted under the foot. From these signals, stride-to-stride measurements of footfall contact times were derived.
  • Gait in Parkinson’s Disease [248]. It includes gait measurements from 93 patients with idiopathic Parkinson’s disease (mean age: 66.3 years; 63% male) and 73 healthy controls (mean age: 66.3 years; 55% male). Vertical ground reaction force data were recorded as participants walked at their usual, self-selected pace for approximately 2 minutes on level ground. Each foot had 8 sensors that measured force over time. The output from each of these 16 sensors was digitized at a rate of 100 samples per second. Additionally, two signals reflecting the summed outputs of the 8 sensors per foot are included.
  • SDUGait [249]. This dataset includes 52 subjects (28 males and 24 females) with an average age of 22. Each subject has 20 sequences, which contain at least 6 fixed walking directions and 2 arbitrary directions, totaling 1,040 sequences. It was collected using the second-generation Kinect (Kinect V2), which not only provides a broader viewing angle but also produces higher-resolution depth images with 25 body joints.
  • Yogev et al. (also termed Ga dataset) [250]. This dataset collects vertical ground reaction force data of subjects by putting 8 sensors on their feet when subjects walk on level ground. It contains 29 PD patients (9 females, 20 males) with an average age of 61.6 years, an average height of 1.67 meters, and an average weight of 73.1 kilograms. 18 healthy controls (8 females, 10 males) are involved with an average age of 57.9 years, an average height of 1.68 meters, and an average weight of 74.2 kilograms.
  • Frenkel-Toledo et al. (also termed Ju dataset) [251]. This dataset collects vertical ground reaction force data by 8 sensors when the subjects walk on a treadmill. It contains 35 PD patients (13 females, 22 males) and 29 healthy controls (11 females, 18 males).
  • Hausdorff et al. (also termed Si dataset) [252]. This dataset collects vertical ground reaction force data when the subjects move at a comfortable place with Rhythmic Auditory Stimulation (RAS). It contains 29 PD patients (13 females, 16 males) and 26 healthy controls (14 females, 12 males).
  • Ga, Ju, Si Groups [251, 252, 250]. This dataset is a combination of the above three small datasets Ga [250], Si [251], and Ju [252], and is commonly used in gait-based PD classification and diagnosis tasks.
All the above datasets are widely used as public benchmarks to evaluate the performance of models. We further discuss the challenges of existing datasets in Sec. 7.2.
6.2 State-of-The-Art Methods and Performance
We utilize the proposed quality evaluation criterion (see Sec. 4.1) to select the most representative state-of-the-art methods from studies with the highest quality:
Representative State-of-The-Art Methods for Gait-Based PD Diagnosis:
  • In [108] (ID = P-15, $S_{Q}=1.00$), a novel hybrid model based on BP neural networks is proposed to classify PD and HC, which outperforms four baseline methods on a public dataset with an average accuracy of 87.9%.
  • [155] (ID = P-17, $S_{Q}=1.00$) introduces a Q-backpropagated Time Delay Neural Network (Q-BTDNN), which harvests the temporal gait data collected from wearable devices, combined with the Q-backpropagation (Q-BP) algorithm for predicting gait disturbances in PD.
  • [157] (ID = P-75, $S_{Q}=1.00$) uses deep CNN to classify gait deterioration in PD patients based on spatio-temporal vGRF signals and achieves 95.5% accuracy in average. The study suggests that the heel strike and body balance are the most indicative gait elements in PD classification.
  • [158] (ID = P-82, $S_{Q}=1.00$) develops a novel NN model to classify PD and HC based on the vGRF dataset collected from a public dataset and achieves an accuracy result of 84.8%.
  • [159] (ID = P-85, $S_{Q}=1.00$) proposes an interval type-2 fuzzy neural network with novel quasi-Levenberg-Marquardt (qLM) learning approach to identify PD patients from the gait cycle based vGRF signals. Based on the 10 clinical features extracted from 14 gait features, the model achieves 97.6% accuracy with a smaller size of parameters and a good interpretability ($i.e.$, rules based on fuzzy logic).
  • In [160] (ID = P-93, $S_{Q}=1.00$), the authors propose a novel architecture PD-ResNet based on ResNet [100] and use it to differentiate PD and HC. Its result outperforms many existing state-of-the-art models with 95.5% accuracy.
  • In [161] (ID = P-97, $S_{Q}=1.00$), a novel model with CNN and RF is devised to perform PD detection tasks from time and frequency features of vGRF, achieving 99.5% accuracy.
  • In [82] (ID = P-106, $S_{Q}=1.00$) proposes a correlative memory neural network (CorrMNN) scheme to extract spatio-temporal features to discriminate NDs’ gait from normal gait based on vGRF, and achieves up to 100% accuracy.
  • In [109] (ID = P-123, $S_{Q}=1.00$) devises a static-dynamic temporal network that consists of one-dimensional, two-dimensional convolutional networks, and attention mechanisms to classify PD patients based on foot force (vGRF time series) and achieves 96.7% accuracy.
  • Johnson et al. (ID = P-125, $S_{Q}=0.88$) addressed early-stage PD monitoring and diagnostic challenges using consumer-grade wearable devices and sensor data from the WATCH-PD study. Data were gathered over a year from early-stage PD patients (n = 82) and age-matched controls (n = 50). Leveraging disease- and behavior-relevant feature engineering with multivariate machine learning, the authors developed a PD screening tool based on a random forest model. Achieving 92.3% accuracy, 90.0% sensitivity, and 100% specificity (AUC = 0.92), this approach underscores the potential of remotely monitored wearable devices for population-wide PD screening and monitoring.
  • The necessity for objective and efficient gait assessment in PD using the Movement Disorder Society Unified Parkinson’s Disease Rating Scale (MDS-UPDRS) is proposed in [164] (ID = P-126, $S_{Q}=1$). To overcome the limitations of existing PD gait skeleton datasets, they collect a sizable dataset including 102 PD patients, 30 healthy older adults, and 16 young adults using multi-view Azure Kinect sensors. In addition, this paper develops a cross-spatiotemporal graph convolution network (CST-GCN) that captures complex gait features by modeling cross-spacetime dependencies between joints to assess age-related effects on PD gait.
Representative State-of-The-Art Method for Gait-Based AD Diagnosis:
  • [76] (ID = A-2, $S_{Q}=0.89$) proposes an innovative cascade neural network approach that combines gait and EEG data for more accurate and efficient classification of AD. The model utilizes attention-based ST-GCN to extract discriminative features from gait data. Experimental results show superior performance compared to other AD diagnosis methods, with the lower body and right upper limb identified as important for early AD diagnosis.
  • CLADSI [235] (ID = A-6, $S_{Q}=0.77$) addresses the challenge of continuous monitoring and stage identification in Alzheimer’s disease (AD) using motion sensor data. Data from 35 AD patients, monitored over a week via accelerometers in a daycare center, was used to develop a method enabling CNN to learn from a continuous stream of sensor data without retaining prior data. This approach allows CNN to self-configure and adapt as new data is received. The CLADSI achieved accuracy rates of 86.94%, 86.48%, and 84.37% for 2, 3, and 4-stage classification tasks, respectively, demonstrating potential for continual learning applications in long-term patient monitoring without human intervention.
Representative State-of-The-Art Method for Gait-Based ALS Diagnosis:
  • In [237] (ID = S-3, $Q_{S}=0.67$) extracts metrics of fluctuation magnitude and fluctuation dynamics from gait time series, and leverages five classic machine learning models (SVM, KNN, NB, LDA, DT) to classify ALS and healthy subjects. Both SVM and KNN achieves the best performance with 89.7% accuracy on the GaitNDD database.
Representative State-of-The-Art Method for Gait-Based HD Diagnosis:
  • In [241] (ID = H-5, $S_{Q}=0.78$) leverages classic machine learning models (SVM, DT, KNN, NB, LDA) to identify HD from the extracted gait features such as right stance interval.
Representative State-of-The-Art Method for Gait-Based MS Diagnosis:
  • The study of [245] (ID = M-4, $S_{Q}=0.89$) devises a novel CNN-based transfer learning framework to identify MS patients from wearable smartphone sensor data and achieves up to 91.1% accuracy. The interpretations of results also suggest that cadence-based measures, gait speed, and ambulation-related signal perturbations are distinct features for classification.
Representative State-of-The-Art Methods for Gait-Based Combinatorial Diagnosis of Different NDs (PD, ALS, HD):
  • In [114] (ID = C-12, $S_{Q}=0.89$) introduces a non-invasive, cost-effective diagnostic tool for analyzing gait signals obtained from vGRF sensors. The best classification results are obtained using SVM model and the achieved average accuracy rates are 100% for ALS, 99.78% for PD, and 99.90% for HD.
  • A hybrid model that combines multiple sensor data is proposed in [82] (ID = C-15, $S_{Q}=0.89$) to learn gait differences between ALS, HD, PD, and healthy individuals. The model includes a spatial feature extractor (SFE) and a correlative memory neural network (CorrMNN) for capturing representative features and temporal information, which outperforms many state-of-the-art techniques with 99.7-100.0% accuracy.
  • A reliable diagnostic model is proposed in [230] (ID = C-20, $S_{Q}=0.77$) for neurodegenerative diseases using gait data transformed into QR codes and classified with CNN. This model, trained on QR-encoded gait data from patients (PD: n = 15, HD: n = 20, ALS: n = 13) and healthy controls (n = 16), achieved high accuracy in distinguishing neurodegenerative conditions: 94.86% for NDD vs. control, 95.81% for PD vs. control, 93.56% for HD vs. control, 97.65% for ALS vs. control, and 84.65% for PD vs. HD vs. ALS vs. control. These results highlight the potential of this approach as a complementary diagnostic tool, especially for individuals exhibiting different levels of motor impairment.
  • A multimodal gait-abnormality-recognition framework named CNN-BiLSTM is proposed in [231] (ID = C-21, $S_{Q}=1$) to enhance neurological disease detection, addressing limitations in single-sensor approaches. This framework processes data from multiple sensors, utilizing an adaptive sliding window and converting time series into time-frequency plots for unified multi-sensor representation. In testing, the framework achieved 98.89% accuracy in Parkinson’s severity classification and high recognition rates for ALS (100%), PD (96.97%), and HD (95.43%). The adaptability and efficiency of CNN-BiLSTM make it promising for daily gait monitoring and personalized rehabilitation planning.
As these studies with the highest quality scores possess higher novelty, more comprehensive comparison, and more sufficient experimental samples that are crucial for creating advanced models, the AI models in these studies can be viewed as state-of-the-art representatives for the diagnosis of different NDs. Notably, both classic machine learning models (6 studies, 30.0%) and deep learning models (14 studies, 70.0%) are valuable for NDs diagnosis tasks and can achieve highly competitive performance. However, there are still many challenges (e.g., generalizability) confronted by these AI models, and we further discuss them in Sec. 7.4.
Representative State-of-The-Art Multi-Modal Approaches for Gait-Based Diagnosis of NDs
  • These two following articles both employed Microsoft Kinect for data collection. First article [84] extracted postural and kinematic features (gait, finger tapping, and foot tapping) for PD classification. A total of nine features for gait analysis were used to train an ANN classifier to perform PD classification with 89.4% accuracy. Second study [85] presented a digital signal processing methods that to extract gait features such as leg length, stride length, and gait velocity. Results showed accurate classification using neural networks, indicating the potential use of Kinect sensors for gait analysis and disorder recognition. The study discussed the possibility of using Kinect as an inexpensive alternative to complex camera systems.
  • This work used a marker-based IR camera and a force-sensitive platform to derive gait parameters such as spatiotemporal, kinematic, and kinetic features. It then utilized two machine learning model, SVM and ANN, to classify PD and HC by using those features and achieved a high accuracy [88].
  • This work [83] combined conventional camera, IMUs, and a force platform for data collection. It also constitutes an experimental protocol for the assessment of PD motor signs of lower extremities, and leverages classic machine learning models to classify PD patients from 18 gait features extracted from gait data and achieved an 88% accuracy.
  • This article [82] utilized not only Kinect, but also vGRF and IMU sensors to collect data and proposed a correlative memory neural network scheme to extract spatio-temporal features and discriminate NDDs’ gait from normal gait and achieved an up to 100% accuracy. The proposed model can also be used to perform gait recognition which outperformed existing state-of-the-art methods.
@@ 7.2 Scarcity and Imbalance of Data
The medical data including gait data collected for NDs diagnosis are highly limited [253], and in the selected 164 articles only several datasets are publicly available: The Gait in Neurodegenerative Disease Database and Gait in Parkinson’s Disease from Physionet database [248] are two most frequently-used used datasets; Other datasets include SDUGait [249], Floodlight PoC dataset [254], UK Biobank Dataset [255], Ga, Ju, Si Groups [251, 252, 250]. However, these datasets are limited in both sample sizes and NDs types. Most of them contain less than 100 subjects (see Sec. 3.2.7), and the number of NDs types is typically less than 3, while PD gait data are much more than other NDs in terms of quantity and quality ($i.e.,$ unbalanced data size of different diseases). This could limit the learning performance of AI models especially deep learning models, which inherently require large-scale and class-balanced data [256] to achieve more reliable prediction.
@@ 7.3 Lack of Multi-Source/Modal Gait Data
Therefore, an important future direction is to collect gait data from multiple sources and modalities [253]. First, it is feasible to combine multiple devices as different sources of same-type data ($e.g.,$ multi-view data with differently-positioned cameras), where all devices/sources can compensate each other to better reconstruct or depict the whole pattern with less information loss [258]. Second, combining different modalities such as RGB images, depth images, 3D skeletons, and force data is valuable as they can provide gait information from different dimensions ($e.g.,$ appearances, poses) to better predict NDs [82].
@@ 7.4 Challenges in AI Model Design
To address the aforementioned challenges, it is essential to improve not only the accuracy of an AI model but also its generalizability, efficiency, and interpretability. A potential future direction is to exploit more advanced deep learning models and larger-scale gait datasets with different scenario settings to learn both domain-specific ($e.g.,$ AD-specific) and domain-general ($i.e.,$ NDs-shared) gait features, so as to improve the model accuracy and enable it to be transferred/generalized to other scenarios. For efficiency, it is feasible to employ lighter network architecture ($e.g.,$ MobileNets [266]) or model compression techniques ($e.g.,$ knowledge distillation [267], network pruning [268]) to reduce the model size and computational complexity but retaining similar model accuracy. Another promising direction is to adopt smaller gait data such as the emerging 3D skeleton data, which can concisely represent human poses with key body joints and typically require lightweight AI models to learn discriminative gait patterns [269]. To improve model interpretability, different types of human-friendly explanation including text description and feature visualization can be considered. We can integrate model-specific explanation mechanisms into different AI architectures, $e.g.,$ visual class activation maps (CAM) [270] for CNN, knowledge graphs [271] for GNN. Inspired by the recent success of ChatGPT, it can be used to provide insightful description for the importance of features if we transform gait features into text format. Moreover, ChatGPT can also act as an agent to lead and improve the AI model learning [272], in which we can add prompts to generate interpretable results. As there exist key body parts and poses for NDs diagnosis, we can try to disentangle the gait abnormality related semantics from normal motion semantics ($e.g.,$ daily actions) to identify the most important or unique component of NDs patients’ features. Simultaneously improving the accuracy, generalizability, efficiency, and interpretability, it is hopefully to obtain a more reliable and practical AI model for gait-based NDs diagnosis. Meanwhile, a unified evaluation protocol, which synergizes different benchmark datasets, validation settings, and evaluation metrics, should be further devised to fairly compare different AI models to facilitate the development of this area.
@@ 8 Conclusion
In this paper, we provide an overview of the recent advancements of AI models for diagnosing NDs through human gait. Our survey showcases a general process of gait-based NDs diagnosis using AI with focus on five representative NDs: PD, AD, HD, ALS, MS, and systematically reviews all existing studies with an elaboration on their used gait data, AI models, and overall performance. A novel quality evaluation criterion is also devised to quantitatively assess the quality of surveyed studies. We outline and discuss the current challenges, potential solutions, and promising future directions in this field. Furthermore, we propose a novel research vision on the utilization of emerging 3D skeleton data for human gait representation and AI model development for more efficient NDs diagnosis.
@@
Acknowledgements
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD/2022-01-034[T]).